OpenStudy (anonymous):

What is the slope of the line 2x - 5y = -10?

OpenStudy (caozeyuan):

-5y=-10-2x, y=2+2/5x so the slope is 2/5

OpenStudy (anonymous):

Start with the given equation Let's find the x-intercept To find the x-intercept, let y=0 and solve for x: Plug in Simplify Divide both sides by 2 Reduce So the x-intercept is (note: the x-intercept will always have a y-coordinate equal to zero) ------------------ Start with the given equation Now let's find the y-intercept To find the y-intercept, let x=0 and solve for y: Plug in Simplify Divide both sides by -5 Reduce